OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. 2. Describe how you would graph this line using the slope-intercept method. Be

OpenStudy (anonymous):

2x + 3y = 1,200 Change the equation into slope-intercept form. 2x + 3y = 1200 [original equation] 3y = 1200 - 2x [subtract 2x on each side] y = 400 - 2x/3 [divide each side by 3] y = -2/3 x + 400 Identify the slope and the y-intercept of the equation. Slope = -2/3 y-intercept = 400 Describe how you would graph this line using the slope-intercept method. The y-intercept is 400, so a graph of this line would intersect the y-axis at 400. Then, for each 2 steps down, 3 must be taken across.